in a rectangle the width is four more than one fifth of the length. The perimeter is 32 inches. Find the length and width of the rectangle
Answer by macston(5194)

Length=L Width=W=1/5L+4
Perimeter=P=2(L+W)=32 in Substitute for W
2(L+(1/5L+4))=32 in
2L+2/5L+8=32 in subtract 8 from each side
2 2/5 L+8-8=32-8 in convert to decimal
2.4L=24 divide both sides by 2.4
2.4L/2.4=24/2.4 in
L=10
CHECK:
If L=10, W=1/5(10)+4=2+4=6
P=2(L+W)=32 in
P=2(10+6)=32 in
P=32in=32in